Neutron stars are some of the densest objects in the universe. They form during supernova ex-


plosions and are very small compared to other astronomical objects. You can assume 5 x 1017


kg/m3 as the average density of a neutron star. The Earth has a total mass of about 5.97 x 1024 kg.


What would be the diameter of the Earth if it had the density of a neutron star?

1
Expert's answer
2022-03-01T15:17:50-0500

"d=\sqrt[3]{\frac{6m}{\pi \rho}}=283.6~m."